### v1

# V2XPnP: Vehicle-to-Everything Spatio-Temporal Fusion for Multi-Agent Perception and Prediction

Abstract
Vehicle-to-everything (V2X) technologies offer a promising paradigm to mitigate the limitations of constrained observability in single-vehicle systems. Prior work primarily focuses on single-frame cooperative perception, which fuses agents’ information across different spatial locations but ignores temporal cues and temporal tasks (e.g., temporal perception and prediction). In this paper, we focus on temporal perception and prediction tasks in V2X scenarios and design one-step and multi-step communication strategies (when to transmit) as well as examine their integration with three fusion strategies - early, late, and intermediate (what to transmit), providing comprehensive benchmarks with various fusion models (how to fuse). Furthermore, we propose V2XPnP, a novel intermediate fusion framework within one-step communication for end-to-end perception and prediction. Our framework employs a unified Transformer-based architecture to effectively model complex spatiotemporal relationships across temporal per-frame, spatial per-agent, and high-definition map. Moreover, we introduce the V2XPnP Sequential Dataset that supports all V2X cooperation modes and addresses the limitations of existing real-world datasets, which are restricted to single-frame or single-mode cooperation. Extensive experiments demonstrate our framework outperforms state-of-the-art methods in both perception and prediction tasks.
footnotetext: Corresponding author {zhiyuh,jiaqima}@ucla.edu
1 Introduction
Autonomous driving systems are required to accurately perceive surrounding road users and predict their future trajectories to ensure safe and interactive driving. Despite recent advances in perception and prediction [49, 22, 16], single-vehicle systems still struggle with limited perception range and occlusion issues [42, 24], compromising driving performance and road safety. Consequently, vehicle-to-everything (V2X) technologies have emerged as a promising paradigm to address these challenges, which enable connected and automated vehicles (CAVs) and infrastructures to share complementary information and mitigate occlusions, thereby supporting holistic environment understanding [35, 18, 12].
Figure 1: Illustration of V2X temporal tasks and our V2X spatio-temporal fusion framework. By incorporating temporal information, our framework enhances V2X communication and supports end-to-end perception and prediction beyond single-frame perception.
Despite their potential, existing works focus on frame-by-frame cooperative detection [41, 37, 51], which aggregates information from agents at different spatial locations. However, these works overlook temporal cues across sequential frames, which are important for locating previously visible but currently undetected objects [44] and predicting objects’ future trajectories [29]. The potential of V2X for enhancing these temporal tasks, particularly in cooperative temporal perception and prediction, remains largely unexplored. We aim to address critical questions in multi-agent cooperation: (1) What information to transmit? (2) When to transmit it? (3) How to fuse information across spatial and temporal dimensions? To address what to transmit, we expand the three fusion strategies in single-frame cooperative perception (i.e., early, late, and intermediate) to incorporate the temporal dimension. Regarding when to transmit within the temporal dimension, we introduce one-step and multi-step communication strategies to capture multi-frame temporal information. For how to fuse, we conduct a systematic analysis across various spatio-temporal fusion strategies, providing comprehensive benchmarks for cooperative perception and prediction tasks across all V2X collaboration modes.
Among these strategies, we advocate intermediate fusion within one-step communication, because it effectively balances the trade-off between accuracy and increased transmission load. Moreover, its capability to transmit intermediate spatial-temporal features makes it well-suited for end-to-end perception and prediction, allowing for feature sharing across multiple tasks and reducing computational demands, as shown in Fig. 1. Based on this strategy, we propose V2XPnP, a V2X spatio-temporal fusion framework leveraging a unified Transformer structure for effective spatial-temporal fusion, encompassing temporal attention, self-spatial attention, multi-agent spatial attention, and map attention. Each agent first extracts its inter-frame and self-spatial features, which can support single-vehicle perception and prediction while reducing the communication load, and then the multi-agent spatial attention model fuses the single-agent feature across different agents.
Another challenge is the lack of real-world sequential datasets encompassing diverse V2X collaboration modes. In V2X scenarios, vehicles and infrastructures serve as primary agents, with collaboration modes that include vehicle-to-vehicle (V2V), vehicle-to-infrastructure (V2I), and infrastructure-to-infrastructure (I2I) [18]. Most existing datasets [38, 45] are non-sequential, limited to single collaboration modes, and focus only on single-frame cooperative perception, lacking support for temporal tasks. To bridge this gap, we introduce the first large-scale real-world V2XPnP Sequential Dataset, featuring four agents and supporting all collaboration modes. This dataset includes temporally consistent perception and trajectory data across 100 vehicle-centric (VC) scenarios and 63 infrastructure-centric (IC) scenarios, totaling 40k frames, along with point-cloud and vector maps from 24 collected intersections. The main contributions of this paper can be summarized as follows:
  1. 1.
    We present V2XPnP, a V2X spatio-temporal fusion framework with a novel intermediate fusion model within one-step communication. This framework is based on unified Transformer architecture integrating diverse attention fusion modules for V2X spatial-temporal information.
  2. 2.
    We introduce the first large-scale, real-world V2X sequential dataset featuring multiple agents and all V2X collaboration modes (i.e., VC, IC, V2V, I2I), encompassing perception data, object trajectories, and map data.
  3. 3.
    We conduct extensive analysis across various spatio-temporal fusion strategies and provide comprehensive benchmarks for cooperative perception and prediction in all V2X collaboration modes, demonstrating the state-of-the-art performance of the proposed model in cooperative temporal perception and prediction tasks.
2 Related Work
Figure 2: Illustration of various V2X fusion strategies for perception and prediction. (a) What to Transmit: early, intermediate, and late fusion, transmitting raw sensor data, intermediate BEV features, or bounding boxes. (b) When to Transmit: multi-step and one-step communication.
End-to-end Perception and Prediction. Safe autonomous driving fundamentally depends on accurate perception and prediction [17, 25, 39, 47]. In single-agent systems, significant efforts have been dedicated to temporal perception and prediction [26, 17], with numerous end-to-end perception and prediction frameworks emerging [29, 27, 11]. These frameworks offer computational efficiency by sharing information across tasks and mitigating error propagation inherent in modular architectures. For instance, FaF [29] employed two strategies to extract the temporal features and directly output the results. PnPNet [27] incorporated a loop tracking module to improve performance. Recently, Bird’s-Eye-View (BEV)-based approaches have gained prominence, focusing on leveraging camera data [11, 14].
V2X Perception and Prediction. Extensive studies have explored V2X perception, with intermediate fusion as a widely adopted approach. OPV2V [42] proposed an attention-based fusion pipeline, while V2X-ViT [41] and CoBEVT [40] utilized vision Transformer to fuse spatial features. However, coordinating a multi-agent system for temporal tasks remains an open challenge. In the prediction domain, extensive exploration of deep learning models, from CNN and LSTM [8] to Transformer [33] and graph network [19], has been conducted to model inter-agent interactions. However, the limited short-term visibility of individual vehicles continues to restrict prediction accuracy. Cooperative prediction leveraging V2X has shown potential, though research remains preliminary [46, 36, 44]. For instance, trajectory-level fusion has demonstrated improved prediction accuracy compared to no fusion [30]. Moreover, V2Vnet [35] utilized a graph neural network to perform joint perception and prediction. Nevertheless, the field currently lacks a comprehensive framework for V2X spatio-temporal tasks, and further research is still needed to fully realize the advantages of perception and prediction offered by V2X systems.
Real-world Driving Datasets. Public driving datasets have been instrumental in advancing autonomous driving research. Early sequential datasets, such as NGSIM [1] and HighD [21], provide object trajectories on highways [50], but lack perception data. The following datasets - nuScenes [4], Argoverse [5], and Waymo [34] - collected real perception data with AVs and extended to urban environments. However, these datasets can not support V2X research due to the single-agent view. To address this, simulated V2X datasets, such as V2XSet [41], were developed. Recently, datasets including Dair-V2X [45], V2V4Real [43], RCooper [13], and V2X-Real [38] contributed significant advancements in real-world data across V2I, V2V, I2I, and V2X perception tasks. However, the lack of real-world sequential V2X datasets remains. The only work in this field, V2X-Seq [46], only provides sequential V2I data with download restrictions.
3 V2XPnP Fusion Framework
The cooperative temporal perception and prediction task requires the integration of temporal information across historical frames and spatial information from multiple agents. It is defined as follows: given map and historical sequences of raw perception data ${\mathbf{P}_{i}^{t}},i\in\{1,\cdots,N\}$, from all $N$ agents within the communication range of the ego agent, the objective is to detect objects in the current frame and predict their future trajectories considering the map information.
3.1 V2X Spatio-Temporal Fusion Strategies
This section addresses what and when to transmit for spatio-temporal fusion, and Sec. 3.2 delves into the fusion process.
What to Transmit. In multi-agent spatial fusion, three fusion strategies are widely adopted in single-frame cooperative perception, i.e., early, late, and intermediate fusion [15, 45], which involve the transmission of raw perception data, bounding boxes, and intermediate features, respectively. Adopting this framework, we extend these fusion approaches to the V2X spatio-temporal fusion context, as illustrated in Fig. 2. (1) Early fusion transmits the entire raw historical perception data to retain complete feature information but imposes the highest transmission load. (2) Late fusion shares only the final detected results at each historical frame, resulting in the lowest transmission load but losing most of the feature information. (3) Intermediate fusion transmits intermediate spatio-temporal BEV features, striking a balance between information quality and transmission load.
When to Transmit. Determining when to transmit in spatio-temporal fusion is more challenging due to the inclusion of temporal data as compared to the existing spatial fusion. A straightforward approach is Multi-step Communication, where each transmission only includes the current frame’s data. However, the multiple steps cause the accumulation of delays and data loss, and obtaining complete historical data requires that neighboring agents stay within the ego agent’s limited communication range throughout historical frames. In practice, the ego agent should aggregate as much data as possible from other agents within a single transmission, rather than relying on multiple exchanges to obtain complete spatio-temporal information. To address this, we propose a One-step Communication strategy, where individual agents share all their historical data within a single communication.
Intermediate Fusion with One-step Communication. Synthesizing the considerations for what to transmit and when to transmit, we adopt an intermediate fusion within the one-step communication strategy, which fuses temporal data from multiple frames before transmission, allowing each agent to share aggregated information without excessive communication overhead. In multi-step intermediate fusion, each agent transmits BEV feature maps at every timestep $\mathbf{F}_{i}^{t}\in\mathbb{R}^{H\times W\times C}$, which denotes agent $i$’s BEV feature at time $t$ with height $H$, width $W$, and channels $C$. The cumulative data shared across $T$ frames is a stacked sequence $\mathbf{F}_{i}^{seq}\in\mathbb{R}^{T\times H\times W\times C}$, resulting in a significantly higher transmission load than single-frame, along with potential delays and data loss. Conversely, in our proposed strategy, each agent first fuses its historical BEV features internally, reducing the sequence from $\mathbf{F}_{i}^{seq}$ to a single condensed BEV feature map $\mathbf{F}^{{}^{\prime}}_{i}\in\mathbb{R}^{H\times W\times C}$. This reduction allows agents to transmit a compact data packet, comparable in size to single-frame cooperative perception, thereby conserving bandwidth while preserving essential spatio-temporal information.
Figure 3: The V2XPnP framework and multi-agent spatio-temporal fusion model. The framework comprises various components for feature extraction, fusion, and decoding. Within our fusion model, we introduce multiple attention mechanisms to enhance the fusion process.
3.2 V2XPnP Framework
The spatio-temporal features of intermediate fusion render it a natural fit for end-to-end perception and prediction. Accordingly, we propose a unified end-to-end perception and prediction framework to perform multiple tasks across spatio-temporal dimensions. The overall V2XPnP framework is illustrated in Fig. 3, which includes six components and is unfolded in this section. The detail of the spatio-temporal fusion model is provided in Sec. 3.3. Notably, each module in V2XPnP is modular, allowing for easy replacement.
V2X Metadata Sharing. Each agent in the V2X system is an observer and collaborator in the shared environment. Each agent will first determine its collaborators with its communication range, and share the metadata, such as relative poses and extrinsic, to construct a spatial V2X graph. Each node and edge in the graph represents an agent and a communication channel. All the point clouds are transformed into the ego agent’s coordinate frame before feature extraction.
LiDAR Feature Extraction. We utilize the PointPillar network [22] to extract the LiDAR feature for each agent $i$ at time $t$, which has low inference latency. The extracted features are structured into a 2D pseudo-image representation to produce salient feature maps $\mathbf{F}_{i}^{t}$, and the sequential feature is stacked as $\mathbf{F}^{seq}_{i}\in\mathbb{R}^{T\times H\times W\times C}$
Multi-frame Temporal Fusion. We propose a Transformer-based temporal fusion module to iteratively perform inter-frame and intra-agent BEV feature fusion through self-attention mechanisms. The spatio-temporal feature of each agent is extracted through the temporal dimension while minimizing communication overhead. The feature map for each agent after temporal fusion is $\mathbf{F}^{{}^{\prime}}_{i}\in\mathbb{R}^{H\times W\times C}$.
Compression and Sharing. To reduce the transmission load, intermediate features are compressed using a $1\times 1$ kernel convolution network in its channel dimension. The ego agent uses another convolution network to decompress the features, restoring them to their original dimensionality.
Multi-agent Spatial Fusion. The decompressed features are passed into a Transformer-based multi-agent spatial fusion network to learn inter-agent and intra-agent spatio-temporal interaction and update the multi-agent feature map $\mathbf{F}^{{}^{\prime}}$.
Map Feature Extraction. The HD map is directly accessed for each agent without V2X fusion. We project the vectorized HD map to BEV space by incorporating the map polylines into each BEV feature grid. We first employ a multi-layer perceptron ($\operatorname{MLP}$) to encode the surrounding map polylines for each grid $\mathbf{M}\in\mathbb{R}^{H\times W\times N_{m}\times n\times D}$, resulting in the map feature $\mathbf{F}_{m}\in\mathbb{R}^{H\times W\times N_{m}\times C}$. Here, $N_{m}$ and $n$ represent the number of map polylines and the waypoints per polyline, while $D$ represents waypoint attributes (i.e., position and lane type). The map encoding is expressed as:
$\mathbf{F}_{m}=\phi\left(\operatorname{MLP}\left(\mathbf{M}\right)\right),$ (1)
where $\phi$ denotes max-pooling on the waypoint axis. Then, a map-BEV attention is introduced to inject the map feature into the BEV feature. We concatenate the BEV and map feature $\mathbf{F}_{bm}=[\mathbf{F}^{{}^{\prime}},\mathbf{F}_{m}]\in\mathbb{R}^{H\times W\times(1+N_{m})\times C}$, and add the position embedding $\mathbf{P}_{m}$ based on sinusoidal positional encoding. The final content feature map $\mathbf{F}\in\mathbb{R}^{H\times W\times C}$ is updated by multi-head self-attention ($\operatorname{MHSA}$) as follows:
$\mathbf{F}\!=\!\operatorname{MHSA}\bigl{(}\operatorname{Q}\colon[\mathbf{F}_{bm},\!\mathbf{P}_{m}],\operatorname{K}\colon[\mathbf{F}_{bm},\!\mathbf{P}_{m}],\operatorname{V}\colon\mathbf{F}_{bm}\bigr{)}.$ (2)
Detection and Prediction Heads. Finally, a detection head and a prediction head are connected to the final feature $\mathbf{F}$ to output states for each predefined anchor box: (1) The detection head contains two 2D convolution layers for bounding box regression and classification. The regression branch outputs $(x,y,z,w,l,h,\theta)$, which represents the bounding box position, size, and yaw. The classification branch outputs the confidence score of each anchor box, determining whether it corresponds to an object or background. (2) The prediction head outputs offset values for each anchor box at each future timestamp using two 2D convolution layers, and the final trajectory is generated by accumulating these offsets.
3.3 Spatio-Temporal Fusion Transformer
In this section, we introduce spatio-temporal fusion with a unified Transformer architecture. The proposed model comprises three blocks: temporal attention, self-spatial attention, and multi-agent spatial attention, as shown in Fig. 3, and two core fusion modules. (1) Multi-frame temporal fusion: Each agent first extracts their spatio-temporal features through iterative temporal and self-spatial attentions. (2) Multi-agent spatial fusion: rich BEV features from multiple agents are acquired via V2X and then fused through iterative multi-agent spatial and self-spatial attentions.
Temporal Attention. This block is designed to capture the inter-frame relationship and aggregate historical BEV features $\mathbf{F}^{seq}_{i}$ across the temporal dimension. The history timestamps are encoded with a learnable embedding, which is added to each BEV feature frame to form $\mathbf{F}^{seq^{\prime}}_{i}$. To preserve temporal cues, this block only fuses temporal features from the same spatial positions across frames, and spatial features are further extracted by the self-spatial Transformer. Temporal fusion is expressed as:
$\mathbf{F}_{i}^{tem}$ $=\operatorname{MHSA}\bigl{(}\operatorname{Q}\colon\operatorname{MLP}(\mathbf{F}^{seq^{\prime}}_{i}),\operatorname{K}\colon\operatorname{MLP}(\mathbf{F}^{seq^{\prime}}_{i}),$
$\quad\quad\quad\quad\>\>\operatorname{V}\colon\operatorname{MLP}(\mathbf{F}^{seq^{\prime}}_{i})\bigr{)}.$ (3)
Self-spatial Attention. To capture the intra-agent spatial BEV interaction, this block employs multi-scale window attention to capture spatial features at different resolutions and ranges. A large window focuses on global features for long-term behavior, and a small window preserves local finer information. Note that this block only fuses spatial features for each agent in a frame without inter-frame and inter-agent fusion. Specifically, we utilize local, intermediate, and global windows $P_{k}\in\{P_{loc},P_{inter},P_{gl}\}$ to partition the feature map through the $H$ and $W$ dimension, generating the self-spatial token $\mathbf{F}^{sp}\in\mathbb{R}^{\frac{H}{P_{k}}\times\frac{W}{P_{k}}\times{P_{k}}^{2}\times C}$. With an additional relative position encoding, $\operatorname{MHSA}$ is operated among ${P_{k}}^{2}$ tokens. The final output is obtained by performing split attention to fuse features from different windows.
Multi-agent Spatial Attention. This block facilitates inter-agent fusion by aggregating BEV feature maps across various spatial positions from multiple agents. Considering the different deployment positions and capabilities of vehicle and infrastructure sensors, the multi-agent spatial Transformer is heterogeneous with individual learnable weights for different interaction pairs (i.e., V-I, V-V, I-V, and I-I). The attention token of $i$ th agent $\mathbf{F}_{i,m}^{sp}$ is modulated by its type $m$ embedding, and is weighted by the relation matrix $\mathbf{W}_{\text{att}}^{(e_{i,j})}$ between edge $e_{i,j}$ during aggregation with agent $j$ of type $n$:
$\mathbf{Q}_{i}^{m}\!=\!\operatorname{MLP}(\mathbf{F}_{i,m}),\mathbf{K}_{j}^{n}\!=\!\operatorname{MLP}(\mathbf{F}_{j,n}),\mathbf{V}_{j}^{n}\!=\!\operatorname{MLP}(\mathbf{F}_{j,n}),$
$\mathbf{F}_{i,m}^{sp}=\sum_{j}\operatorname{Softmax}(\mathbf{Q}_{i}^{m}\cdot\mathbf{W}_{\text{att}}^{(e_{i,j})}\cdot\mathbf{K}_{j}^{n})\cdot\mathbf{V}_{j}^{n}.$ (4)
3.4 Learning Objective
The learning objective is comprised of temporal perception and prediction tasks. First, we define the perception loss as the combination of regression $\mathcal{L}_{\text{reg}}$ and classification loss $\mathcal{L}_{\text{cla}}$ of the predefined anchor box. Specifically, the smooth $\ell_{1}$-loss is leveraged for the regression part, and the focal loss [28] is utilized for classification. Second, we define the prediction loss $\mathcal{L}_{\text{pred }}$ as $\ell_{2}$-loss between the prediction points sequence with ground truth trajectory. The final loss function is their weighted sum:
$\mathcal{L}=w_{\text{reg }}\mathcal{L}_{\text{reg }}+w_{\text{cla}}\mathcal{L}_{\text{cla }}+w_{\text{pred }}\mathcal{L}_{\text{pred }},$ (5)
where $w_{\text{reg }}$, $w_{\text{cla}}$, and $w_{\text{pred }}$ represent balance weights.
Figure 4: Illustration of the V2XPnP Sequential Dataset. (a) V2X data acquisition systems; (b) Sequential LiDAR and camera data; (c) Data collection area and vector map; (d) Total number and average tracking length of 3D tracked objects per category.
4 V2XPnP Sequential Dataset
We introduce the V2XPnP-Sequential dataset, the first large-scale, real-world V2X sequential dataset featuring multiple agents and all collaboration modes. This dataset comprises 100 scenarios, each spanning 95 to 283 frames captured at 10 Hz. The dataset comprises two data sequences from CAV perception (point clouds and camera images) and two data sequences from infrastructure perception, as shown in Fig. 4(b). We also provide corresponding vector maps and point-cloud maps for all collection areas, as shown in Fig. 4(c). Ten object categories are included, and the average trajectory length and frequency of each category are shown in Fig. 4(d). Further details on data annotation, trajectory and map generation are provided in the supplementary.
4.1 Data Acquisition
V2X temporal tasks require diverse time-consistent perception data and object behavior data. We choose urban arterial roads, expressways, and intersections to collect V2X data. The sensor configurations for the two CAVs and two infrastructures are shown in Fig. 4(a). By permuting and combining the behavior patterns of CAVs (such as overtaking, platooning, turning, etc.), we designed a total of about 60 interaction pairs between CAVs. From more than 66h driving logs, we annotate 32 representative scenarios and process data from the non-sequential V2X-Real [38] with a V2X sequential data processing pipeline, detailed in Sec. 4.2. Since perception tasks do not need to consider the continuity of object labeling, identical objects in V2X-Real may be assigned different IDs within a sequence, leading to data fragmentation and limiting support for temporal tasks. The final dataset is yielded by processing our collected data and V2X-Real data.
4.2 Sequential Data Processing
Time-consistent data is crucial for temporal tasks, and thus we develop a V2X sequential data processing pipeline to track objects across time and different agents’ views. We construct a multi-agent spatio-temporal graph, where each node $n_{it}^{k}$ represents an annotated bounding box $i$ at time $t$ from agent $k$’s LiDAR data, and edges connect nodes that correspond to the same object. Identifying objects with the same ID then reduces to finding connected components in this graph, with each isolated component representing a unique object. Moreover, we build this graph by leveraging the temporal continuity in single-agent annotations and incorporating multi-agent data. For temporal consistency within a single agent, we add edges between $n_{it}^{k}$ and $n_{it+1}^{k}$, although annotation errors may sometimes assign different IDs to the same object. To address this, we integrate annotations from multiple agents, transforming annotations into a global coordinate frame and connecting nodes if their Intersection-over-Union (IoU) exceeds a threshold. Once the graph is complete, we identify connected components and assign each a unique tracking ID. To mitigate annotation biases, we refine object attributes based on their consensus.
5 Experiments
5.1 Experimental Setup
Evaluation Metrics. Following the detection evaluation protocol in [38], we measure detection performance using Average Precision (AP) at an IoU threshold of 0.5. For prediction, we provide the results of commonly used metrics [32, 10, 31, 3], including Average Displacement Error (ADE), Final Displacement Error (FDE), and Miss Rate (MR) within a 2-meter threshold. However, prediction accuracy is influenced by false positives and missed detections from the perception module. For example, a poorly performing perception module that detects only a simple object in a straight-line trajectory can misleadingly inflate the prediction accuracy. To address this, we employ the End-to-end Perception and Prediction Accuracy (EPA) metric [11] to jointly evaluate perception and prediction performance.
$EPA=\frac{|\hat{S}_{\text{match, hit}}|-\alpha N_{\text{FP}}}{N_{\text{GT}}},$ (6)
where $|\hat{S}_{\text{match, hit}}|$ is the number of true positive objects with prediction $FDE<\tau_{EPA}$, $N_{FP},N_{GT}$ represent the number of false positive objects and ground truth objects, respectively, and $\alpha$ is a penalty coefficient. A higher EPA value indicates superior object detection and prediction capabilities, and we set $\tau_{EPA}=2$m, $\alpha=0.5$ following [11].
Table 1: Benchmark results of cooperative perception and prediction models on V2XPnP Sequential (V2XPnP-Seq) Dataset
Dataset Method E2E Map AP@0.5 (%) $\uparrow$ ADE (m) $\downarrow$ FDE (m) $\downarrow$ MR (%) $\downarrow$ EPA (%) $\uparrow$
V2XPnP-Seq-VC (with V+2I at most) No Fusion 43.9 1.87 3.24 33.8 24.3
No Fusion-FaF* [29] 53.4 1.55 2.81 34.3 31.6
Late Fusion 58.1 1.59 2.82 32.4 33.0
Early Fusion 60.3 1.37 2.49 33.8 36.7
V2VNet* [35] 48.6 2.10 3.75 42.3 25.3
F-Cooper* [6] 66.0 1.35 2.56 36.1 38.7
V2XPnP (Ours) 71.6 +5.6 1.35 2.36 31.7 48.2 +9.5
V2XPnP-Seq-IC (with 2V+I at most) No Fusion 46.4 1.69 3.06 36.2 28.8
No Fusion-FaF* [29] 56.7 1.34 2.65 41.4 31.7
Late Fusion 55.9 1.39 2.44 30.1 32.9
Early Fusion 60.5 1.39 2.63 32.8 39.5
V2VNet* [35] 33.6 1.95 3.53 44.2 16.3
F-Cooper* [6] 60.2 1.21 2.32 36.3 36.3
V2XPnP (Ours) 71.0 +10.8 1.18 2.16 34.0 46.0 +9.7
V2XPnP-Seq-V2V No Fusion 40.8 1.99 3.38 34.0 19.8
No Fusion-FaF* [29] 51.9 1.67 3.12 39.3 27.5
Late Fusion 55.3 1.75 3.07 34.0 30.5
Early Fusion 53.0 1.64 3.11 40.2 26.9
V2VNet* [35] 43.1 3.10 5.55 46.8 19.4
F-Cooper* [6] 60.2 1.69 3.22 41.1 34.4
V2XPnP (Ours) 70.5 +10.3 1.78 3.28 39.9 40.6 +6.2
V2XPnP-Seq-I2I No Fusion 51.0 1.69 3.06 36.2 31.7
No Fusion-FaF* [29] 56.6 1.34 2.65 41.4 31.7
Late Fusion 61.3 1.41 2.50 30.0 41.6
Early Fusion 64.6 1.57 2.98 39.9 37.7
V2VNet* [35] 41.1 1.83 3.34 40.4 23.2
F-Cooper* [6] 58.6 1.34 2.58 40.0 33.6
V2XPnP (Ours) 69.2 +10.6 1.26 2.31 36.5 42.8 +9.2
Collaboration Modes. The V2XPnP Sequential Dataset supports various V2X collaboration modes by organizing data based on specific interaction patterns. Vehicle-Centric (VC): The ego CAV is the focal agent, communicating with other CAVs and infrastructure. Infrastructure-Centric (IC): Infrastructure is the central entity, communicating with other infrastructures and CAVs. Vehicle-to-Vehicle (V2V): The ego CAV communicates exclusively with other CAVs without involving infrastructure. Infrastructure-to-Infrastructure (I2I): Infrastructure units share data only with each other.
Implementation Details. During the testing stage, we select a fixed agent as the ego agent in each cooperative scenario, while the ego agent is shuffled and randomly selected during training. Following the real-time testing and evaluation [48], we set the communication range for the ego agent to 50 meters and evaluate surrounding agents within a range of $x\in[-70,70]$m and $y\in[-40,40]$m. Messages beyond 50 meters are discarded by the ego agent. Besides, the history length is 2s (2Hz), and the prediction horizon is 3s (2Hz). The train/validation/test data splits are 80/6/14 scenarios. All models are trained using the Adam optimizer [20] with early stopping on NVIDIA L40S GPUs. Additional training and model details are provided in the supplementary materials.
5.2 Benchmark Methods
End-to-end methods. We focus on the end-to-end models in this paper and implement an end-to-end model for cooperative temporal perception and prediction. This model uses the same LiDAR backbone and decoding heads with V2XPnP but utilizes a simple temporal fusion module FaF [29] - alternating 2D and 3D convolutions- as the No fusion-FaF baseline. Then, the Early fusion configurations and two state-of-the-art single-frame intermediate fusion models, V2VNet [35] and F-Cooper [6] to integrate the end-to-end model as the early fusion and intermediate fusion baselines. These models marked with are reimplemented in our framework to ensure consistency in their LiDAR backbone, temporal fusion network, and decoding heads.
Decoupled methods. Transmitting final detection results rather than spatio-temporal features renders late fusion incompatible with end-to-end models. Therefore, we benchmark Late Fusion with a decoupled pipeline, where objects in each historical frame are detected using a single-frame perception module, and results are fused via non-maximum suppression [42]. Assuming an ideal tracker to generate object trajectories from perception results and interpolate missing points, we implement an attention-based predictor for trajectory-level prediction tasks, following the prediction model mainstream [33]. To further assess end-to-end performance, we also evaluate a decoupled No-Fusion model.
5.3 Results
Tab. 1 presents the benchmark results across four V2X collaboration modes. Since prediction performance is inherently tied to detection quality, the difficulty posed by different detected objects can significantly impact prediction accuracy. As a result, the EPA metric serves as the most appropriate indicator for assessing the overall performance across models. More details and baseline results are provided in the supplementary material.
Transmit. In both end-to-end and decoupled PnP frameworks, the cooperation perception and prediction performance are consistently better than non-fusion models in all collaboration modes, especially in the primary EPA metric, which demonstrates the benefits of cooperation in temporal perception and prediction. Moreover, intermediate fusion models (e.g., F-Cooper and V2XPnP) generally outperform other fusion strategies, while early fusion consistently surpasses late fusion. However, V2VNet has a worse performance, likely due to the absence of a map and the loss of temporal features during explicit feature ROI matching. Our V2XPnP model archives the best performance, with an average improvement of 15% and 24% over F-Cooper in AP and EPA, respectively, because the unified Transformer model better captures the complex relationship among the spatio-temporal information shared by V2X.
Figure 5: Qualitative results of different fusion models on the testing set. V2XPnP shows better perception and prediction results.
When to Transmit. Tab. 2 shows the performance of our model under two communication strategies. (1) The perception and prediction performance of one-step communication is improved compared to multi-step communication without delay by 5.0% AP and 12% in EPA. This improvement arises because the ego agent directly fuses lossless raw data, avoiding error accumulation from lossy intermediate information transformed across temporal dimensions. (2) The one-step strategy compresses spatio-temporal feature transmission (under a 32$\times$ compression rate) from 5$\times$0.269 Mb to 0.269 Mb compared to multi-step communication, while mitigating information loss when agents move out of communication range in historical frames. At a typical C-V2X data transmission rate [2], the transmission delay of one-step communication is approximately 10$\sim$20 ms. (3) Under the noise settings from [41], we provide the results with 100 ms time delay and positional/head Gaussian noise (0.2m, 0.2°). Due to error accumulation during multiple transmissions, the performance of the multi-step communication strategy decreases by 14%, while the one-step strategy’s performance only decreases by 3.7%, demonstrating greater robustness.
Table 2: Comparison of one-step and multi-step communication; Bold values indicate results under delay and noisy settings.
Strategy AP@0.5 $\uparrow$ ADE $\downarrow$ FDE $\downarrow$ MR $\downarrow$ EPA $\uparrow$
Multi -step 68.2 1.56 2.84 31.8 43.0
56.6 -17% 1.52 2.71 28.0 37.0 -14%
One -step 71.6 1.35 2.36 31.7 48.2
68.5 -4.3% 1.37 2.38 31.6 46.4 -3.7%
How to Fuse. Tab. 3 provides the ablation study of V2XPnP, showing the effectiveness of different components in V2XPnP. The temporal fusion module provides the history information for current-frame detection, while multi-agent spatial fusion alleviates occlusions and improves performance by incorporating other views. The map fusion module enhances trajectory prediction by guiding future trajectories to align with road structures. Our complete V2XPnP model with all these fusion modules performs the best.
End-to-end vs. Decoupled Frameworks. The end-to-end model consistently outperforms the decoupled framework. In no-fusion situations, the end-to-end FaF model outperforms the decoupled model in detection by leveraging temporal cues. Furthermore, FaF achieves performance comparable to late fusion with V2X spatial aggregation due to integrating temporal features. The intermediate fusion of spatio-temporal features in V2XPnP aligns well with the end-to-end architecture, showcasing its superior performance.
Infrastructure vs. Vehicle Centric. As shown in Tab. 1, models under VC and IC modes outperform those in V2V and I2I modes, because VC and IC can aggregate information from up to four agents rather than only two, resulting in enhanced environmental understanding. Additionally, stationary infrastructure-based agents in IC and I2I modes offer higher prediction accuracy by providing elevated sensing perspectives and less noisy sensor data.
Qualitative Results. Fig. 5 visualizes the outcomes of cooperative perception and prediction across different fusion models. The No Fusion model is constrained by its limited field of view and thereby shows subpar performance. FaF*, leveraging temporal information within an end-to-end pipeline, performs better under occlusion. Late and early fusion models significantly benefit from multi-agent data integration, though late fusion remains impacted by error propagation from perception, such as heading errors misleading trajectory direction. Notably, the end-to-end intermediate fusion model, particularly V2XPnP, achieves superior performance in both detection and prediction tasks.
Table 3: Ablation results of V2XPnP model
Temp Spatial Map AP@0.5 $\uparrow$ ADE $\downarrow$ FDE $\downarrow$ MR $\downarrow$ EPA $\uparrow$
43.9 - - - -
57.2 1.52 2.76 35.5 33.8
71.3 1.48 2.70 36.2 44.4
71.6 1.35 2.36 31.7 48.2
6 Conclusions
We propose V2XPnP, a novel V2X spatio-temporal fusion framework for cooperative temporal perception and prediction. The core of this framework is a unified Transformer-based model for spatio-temporal fusion and map fusion. Furthermore, we examine various fusion strategies concerning what, when to transmit, and how to fuse, offering comprehensive benchmarks. Additionally, we introduce the V2X Sequential Dataset, which supports all V2X collaboration modes. Extensive experimental evaluations demonstrate the superior performance of the proposed framework, establishing its efficacy in advancing V2X-enabled cooperative temporal tasks. Future work will focus on developing more effective fusion models and adaptive communication strategies that dynamically optimize bandwidth utilization.
Supplementary Material
A Implementation Details
In this section, we provide detailed configurations for cooperative perception and prediction tasks, including the baseline models used in our experiments and the proposed V2XPnP framework.
A.1 Benchmark Model Details
PointPillar Backbone. For all experiments, we employ the anchor-based PointPillar model [22] as the LiDAR Feature Extraction backbone. The voxel resolution is set to 0.4 meters in both the $x$ and $y$ directions, with a maximum of 32 points per voxel and a total of 32,000 voxels. Additionally, we set the number of anchors per grid cell to 2.
Intermediate Fusion Methods. We implement two state-of-the-art single-frame intermediate fusion methods (i.e., V2VNet [35] and F-Cooper [6]) and extend them to spatio-temporal fusion using a baseline temporal fusion model that alternates between 2D and 3D convolutions, denoted as V2VNet and F-Cooper. The model setting and configurations in multi-agent fusion follow the original implementation in [35] and [6]. Specifically, V2VNet employs direct max-pooling, while F-Cooper utilizes spatial-aware message-passing mechanisms to aggregate features across multiple agents. The feature map dimensions in our experiments are $176\times 100$ with $256$ channels.
Map Feature Extraction. HD maps are represented as sets of polylines, with each polyline comprising 10 points. Because the map is projected onto the BEV space, each grid only contains the five nearest polylines. Each waypoint in a polyline contains seven attributes: $(x,y,d_{x},d_{y},type,x_{pre},y_{pre})$, representing position, direction, lane type, and previous position. These attributes are encoded using MLP layers into a $256$ hidden dimension feature, followed by 1 or 2 Transformer layers with two attention heads to model interactions among map elements.
Decoupled Attention Predictor. For the decoupled perception and prediction pipeline, we implement an attention-based predictor for trajectory-level prediction tasks. This predictor utilizes a 1D Convolution + LSTM Network [8] to encode temporal historical trajectories and a Transformer layer to capture the interaction among objects and the map, then an LSTM-based decoder generates the future predicted trajectories. All trajectory data, including historical and predicted trajectories, are represented in the local coordinate frame of each object.
A.2 V2XPnP Model Details
Temporal Attention. To capture the temporal dependence, we initialize the historical timestamp sequence using Sinusoidal positional encodings conditioned on time and further process these encodings through a Linear layer. The temporal attention block in the multi-frame temporal fusion module has four attention heads. To enhance the inter-frame feature representation, we stack three temporal fusion modules with the temporal attention block.
Self-spatial Attention. This block is applied following either the temporal attention or the multi-agent spatial attention. In self-spatial attention, the feature map is partitioned into patches using common window sizes of $(2,4,8)$. Given the complexity of spatio-temporal fusion across multiple agents, the self-spatial attention module employs a higher number of attention heads $(16,8,4)$ after multi-agent spatial fusion, compared to the heads $(8,4,2)$ used following temporal attention.
Multi-agent Spatial Attention. Our dataset categorizes agents as infrastructure agents, denoted by negative labels (i.e., $-1$ and $-2$), or connected automated vehicles (CAV) agents, denoted by positive labels (i.e., $1$ and $2$). To capture the heterogeneous dependencies among these agents, we construct a heterogeneous graph and employ distinct attention fusion parameters for each agent type. The multi-agent spatial attention utilizes eight attention heads, and we stack three multi-agent spatial fusion modules with the multi-agent spatial attention to capture the inter-agent relationships.
A.3 Loss Function
This section provides the loss function employed in our multi-task model. The initial weights of regression loss $\mathcal{L}_{\text{reg }}$, classification loss $\mathcal{L}_{\text{cla}}$ and prediction loss $\mathcal{L}_{\text{pred}}$ are set as $w_{\text{reg}},w_{\text{cla}},w_{\text{pred}}=2.0,1.0,2.0$. For single-task learning, the same loss function is used but weights exclusively on the components relevant to that task.
Perception Loss. The perception task loss combines classification and regression components, designed to align predicted anchor boxes with ground truth labels. For classification, which involves identifying objects and background elements, we employ Focal Loss [28] to address the imbalance between foreground and background samples. The Focal Loss is expressed as:
$\mathcal{L}_{\text{cla}}=-\alpha(1-p_{t})^{\gamma}\log(p_{t}),$ (S1)
where $p_{t}$ is the predicted probability for the target anchor box, and $\alpha$ and $\gamma$ are balancing and focusing factors. Anchor-wise weights are applied to further enhance the balance between positive and negative samples.
For the regression component, we employ Smooth $\ell_{1}$-Loss to optimize the predicted bounding boxes to match the ground truth labels in terms of position and orientation, and a sine-cosine encoding is employed to handle rotational ambiguities. The Smooth $\ell_{1}$-Loss is defined as:
$\mathcal{L}_{\text{reg}}=\begin{cases}0.5\cdot\frac{\Delta^{2}}{\beta},&\text{if}|\Delta|<\beta,\\ |\Delta|-0.5\cdot\beta,&\text{otherwise},\end{cases}$ (S2)
where $\Delta=\text{prediction}-\text{target}$, and $\beta$ is a hyper-parameter controlling the transition between $\ell_{1}$ and $\ell_{2}$ loss.
Prediction Loss. We adopt the $\ell_{2}$-loss function to minimize the discrepancy between the predicted trajectory and the ground truth.
$\mathcal{L}_{\text{pred}}=\frac{1}{N_{\text{det}}}\frac{1}{T_{\text{valid}}}\sum_{i=1}^{N_{\text{det}}}\sum_{t=1}^{T_{\text{valid}}}\|\boldsymbol{\mu}^{i}_{t}-\mathbf{x}^{i}_{t}\|^{2},$ (S3)
where $\boldsymbol{\mu}^{i}_{t}$ and $\mathbf{x}^{i}_{t}$ represent the predicted position and target position of the $i$-th object at time step $t$. ${T_{\text{valid}}}$ is the number of valid future time steps for the agent, and $N_{\text{det}}$ is the number of detected objects.
A.4 Training Strategy
The end-to-end cooperative perception and prediction model addresses two distinct yet interrelated tasks while integrating information across both temporal and spatial dimensions. Training such an end-to-end model from scratch often results in suboptimal performance, due to the inherent complexity of jointly optimizing these tasks and dimensions. To effectively handle these challenges, we adopt a multi-stage training strategy to progressively refine the model’s capabilities.
Multi-Stage Training Strategy. Initially, the end-to-end perception and prediction model is trained in a single-agent setting, focusing on temporal fusion without incorporating multi-agent spatial fusion. It simplifies the optimization process, enabling the model to learn robust temporal features in isolation. The resulting single-agent model then serves as a pre-trained model for subsequent multi-agent spatial fusion training in the V2X environment. This staged training strategy ensures that the model incrementally acquires the ability to handle the complexities of cooperative perception and prediction tasks.
Stage 1: Single-Agent Multi-task Learning. The single-agent model training stage addresses the core challenge of coordinating multi-task learning to capture complex patterns across perception and prediction tasks. Prediction task requires a comprehensive understanding of objects’ temporal information and their intricate motion patterns, while detection focuses mainly on identifying objects in the current frame, with historical information providing supplementary context. Training both tasks jointly without proper initialization risks overfitting to simpler current-frame features, thereby neglecting the rich but complex temporal features essential for accurate prediction. Moreover, perception is foundational to prediction, as detecting an object is a prerequisite for predicting its motion. To effectively balance the two tasks, we adopt a task-specific training strategy. (1) Single-Frame Perception Training: the training begins by optimizing the model for single-frame perception, establishing a foundation for object detection. (2) Temporal Prediction Training: the prediction task is introduced by freezing the parameters of the detection backbone and training an additional temporal network and prediction head, guiding the model to focus more on the prediction task and effectively learn complex temporal dependencies. (3) Joint Fine-Tuning: the entire model is unfrozen, enabling end-to-end fine-tuning across both tasks.
Table S1: Benchmark results for cooperative temporal perception. No Temp: single-frame perception, FaF: temporal perception with alternating 2D and 3D convolutions, V2XPnP: temporal perception with temporal attention modules.
Dataset No Fusion (AP@0.5 (%) $\uparrow$) Early Fusion (AP@0.5 (%) $\uparrow$) Intermediate Fusion (AP@0.5 (%) $\uparrow$)
No Temp FaF V2XPnP No Temp FaF V2XPnP No Temp FaF V2XPnP
V2XPnP-Seq-VC 43.9 57.1 60.3 63.5 67.0 71.0 65.1 70.3 74.0
V2XPnP-Seq-IC 46.4 61.1 64.7 61.0 65.5 71.4 61.1 67.1 73.2
V2XPnP-Seq-V2V 40.8 53.7 59.1 54.9 56.4 66.6 58.0 61.4 69.4
V2XPnP-Seq-I2I 51.0 61.2 64.7 63.4 66.0 71.6 58.5 62.9 72.4
Stage 2: Multi-Agent Spatio-Temporal Learning. Based on the pre-trained single-agent model, the multi-agent fusion module is introduced and jointly trained with the entire model. At this stage, the primary focus is to balance the two tasks, ensuring that neither perception nor prediction dominates the training process. To achieve this, we employ a dynamic loss-weighting strategy that gradually increases the weight assigned to the prediction loss. This approach ensures balanced optimization, avoiding performance trade-offs between tasks and improving overall effectiveness across both perception and prediction objectives.
Training Details. The model is trained using the Adam optimizer with an initial learning rate of $2\times 10^{-3}$ and a weight decay of $1\times 10^{-4}$. We employ $4$ training stages detailed before, and each training stage consists of $20$ epochs with a batch size of $2$, and early stopping is employed to prevent overfitting. We carefully tune the hyperparameters to ensure the stability and efficiency of the training process.
Table S2: Benchmark results for traditional prediction. No Fusion: prediction based on the no-fusion perception results. Late Fusion: prediction based on the late fusion perception results. Ground Truth: prediction based on the ground truth trajectories with no occlusion or perception errors.
Dataset Method Attention Predictor LSTM Predictor
AP@0.5(%) $\uparrow$ ADE(m) $\downarrow$ FDE(m) $\downarrow$ MR(%) $\downarrow$ AP@0.5(%) $\uparrow$ ADE(m) $\downarrow$ FDE(m) $\downarrow$ MR(%) $\downarrow$
V2XPnP-Seq -VC No Fusion 43.9 1.87 3.24 33.8 43.9 2.91 4.77 35.0
Late Fusion 58.1 1.59 2.81 34.3 58.1 2.76 4.60 33.7
Ground Truth - 0.60 1.26 23.0 - 0.66 1.31 23.0
V2XPnP-Seq -IC No Fusion 46.4 2.10 3.75 42.3 46.4 2.11 3.67 35.8
Late Fusion 55.9 1.39 2.44 30.1 55.9 2.61 4.40 32.7
Ground Truth - 0.63 1.35 26.2 - 0.61 1.31 25.0
V2XPnP-Seq -V2V No Fusion 40.8 1.99 3.38 34.0 40.8 2.98 4.82 34.4
Late Fusion 55.3 1.75 3.07 34.0 55.3 2.87 4.79 35.0
Ground Truth - 0.60 1.26 22.9 - 0.66 1.31 22.8
V2XPnP-Seq -I2I No Fusion 51.0 1.69 3.06 36.2 51.0 2.11 3.67 35.9
Late Fusion 61.3 1.41 2.50 30.0 61.3 2.44 4.18 32.1
Ground Truth - 0.63 1.35 26.2 - 0.61 1.31 25.0
B Cooperative Temporal Perception Task
In addition to the end-to-end perception and prediction task, the sequential nature of our V2XPnP-Sequential dataset facilitates other temporal tasks, including temporal perception and traditional prediction tasks. In this section, we introduce the cooperative temporal perception task and present benchmark results on the V2XPnP-Sequential dataset. Details on the traditional prediction task are provided in Sec. C.
B.1 Problem Formulation
The cooperative temporal perception task is an extension of the single-frame perception task by incorporating historical context. Specifically, given historical $T$ frames raw perception data ${\mathbf{P}_{i}^{t}},i\in\{1,\cdots,N\}$ from all $N$ agents within the communication range of the ego agent, the objective is to detect the surrounding objects in the current frame. The core challenge lies in effectively leveraging temporal information from $T$ past frames to enhance detection accuracy in the present frame.
B.2 Benchmark Methods
For benchmarking, we adapt our end-to-end model, V2XPnP, by removing the prediction head, resulting in a model only for temporal perception. Various V2X fusion strategies are evaluated in this framework, as detailed in Tab. S1. Moreover, we provide another baseline FaF, which adopts a combination of 2D and 3D convolutions for temporal fusion. FaF further integrates with the F-Cooper intermediate fusion method and early fusion method for V2X fusion comparison. We also provide the results of the No Temp model, which excludes temporal fusion and is evaluated using both F-Cooper and early fusion methods. Model parameters and experimental setups for this task are consistent with those used for the end-to-end cooperative perception and prediction task.
B.3 Benchmark Results
The results demonstrate that incorporating temporal cues significantly improves perception performance across all multi-agent fusion strategies. Notably, our V2XPnP model achieves superior results compared to other baselines, due to the careful design of temporal attention. However, we observe a slight performance drop when the same model is applied to the end-to-end cooperative perception and prediction task, compared to its use solely for temporal perception. The possible reason is the difficulty of optimizing both tasks to achieve optimal performance. Nevertheless, the end-to-end model still outperforms other baselines in both perception and prediction tasks. Future research should focus on optimizing the balance between multiple tasks to further enhance the performance of end-to-end models.
C Traditional Cooperative Prediction Task
C.1 Problem Formulation
V2XPnP sequential dataset also supports the traditional prediction task. Compared to end-to-end models, which directly infer future states of objects from perception data, the traditional prediction task forecasts their future trajectories from historical trajectories. The cooperative prediction task is formulated as: given the map and the historical trajectories of all detected objects obtained from the ego agent and other agents (e.g., CAVs and infrastructure units) within the communication range of the ego agent, the objective is to predict future trajectories of these detected objects.
C.2 Benchmark Methods
To investigate the influence of perception results on prediction tasks, we provide two types of input for the prediction models: 1) Ground-truth historical trajectories of surrounding objects; 2) Perception-based historical trajectories generated by the upstream perception module. The first one is the common setting for the traditional trajectory prediction task, assuming full availability of accurate historical trajectories for prediction. However, it ignores real-world challenges such as occlusions and cumulative errors introduced by separate modules. To address this limitation and enable a more realistic evaluation, we designed the second setting, where CAVs can only derive the historical trajectories from the perception results and thus the perception uncertainty can propagate to the downstream prediction. Notably, regardless of the input type, the prediction model is trained using the complete future trajectory dataset aggregated from all agents.
In our experiment, the prediction model configuration and experimental setup align closely with the decoupled attention predictor. Following the LSTM baseline setting in the Waymo motion dataset [9], the LSTM model also serves as a strong baseline, which includes an LSTM encoder and LSTM decoder. We report benchmark results under three configurations: No Fusion, where no perception information is fused; Ground Truth, assuming perfect historical trajectories; Late Fusion, where the decoupled pipeline from the traditional prediction task is employed.
C.3 Benchmark Results
The experimental results, summarized in Tab. S2, compare traditional prediction under three input settings: ground truth trajectories, perception without fusion, and perception with late fusion. The results indicate that as perception improves—from no fusion to late fusion—the prediction performance correspondingly increases. When the environment is fully observable, the task simplifies to the traditional prediction setup, achieving the best overall performance for both detection and prediction. A significant drop in performance is observed for perception-based prediction, highlighting the critical dependency of predictive tasks on perception accuracy. Moreover, the Attention predictor shows better robustness compared to the LSTM baseline under noisy perception inputs, thanks to the attention module for complex interaction feature capturing. We anticipate that this temporal prediction task will inspire further exploration of perception-based prediction approaches.
Figure S1: Examples of interaction scenarios from the V2XPnP-Sequential dataset. The dataset multi-agent perception perspectives and captures diverse interaction behaviors among ten object classes in dense urban traffic environments.
Table S3: Comparison between the V2XPnP-Sequential dataset and other public available driving datasets
Dataset Year Type V2V V2I I2I Trajectory Map
Agent
Number
Tracked
Objects/Scene
3D
Boxes
RGB
Images
LiDAR
Frames
Categories
nuScenes [4] 2019 Real 1 75.75 1.4M 1.4M 400k 23
Waymo Open [9] 2019 Real 1 - 12M 1M 200k 4
OPV2V [42] 2022 Sim 2.89 26.5 230k 44k 11k 1
V2X-Sim [24] 2022 Sim 10 - 26.6k 0 10k 1
V2XSet [41] 2022 Sim 2-7 - 230k 44k 11k 1
DAIR-V2X [45] 2022 Real 2 0 464k 39k 39k 10
V2V4Real [43] 2023 Real 2 - 240k 40k 20k 5
V2X-Seq [46] 2023 Real 2 110 464k 71k - 10
RCooper [13] 2024 Real 2 - - 50k 30k 10
V2X-Real [38] 2024 Real 4 0 1.2M 171K 33k 10
V2XPnP-Seq 2024 Real 4 136 1.45M 208k 40K 10
D V2XPnP Sequential Dataset Details
D.1 Dataset Visualization
Our V2XPnP-Sequential dataset provides two sensor sequences (LiDAR and camera) collected in dense urban environments, capturing diverse interactive behaviors over time. Fig. S1 illustrates two representative interaction scenarios in our dataset, presenting LiDAR and camera data from multiple agents at two key timestamps. The main intersection objects pair have been annotated with red and yellow blocks in different agents’ views.
D.2 Data Acquisition
Sensor Specifications. The dataset was collected using four agents - two connected automated vehicles and two smart infrastructure units. Each CAV is equipped with a RoboSense 128-beam LiDAR, four stereo RGB cameras with 1920 × 1080 resolution, and an integrated GPS/IMU system. The four stereo cameras are mounted on the front, rear, left, and right sides of the CAV, providing a complete 360-degree field of view. Similarly, each infrastructure unit is configured with a 128- or 64-beam LiDAR, two Axis cameras with 1920 × 1080 resolution, and a GPS module. The sensor deployment of our data collection system is shown in Fig. 4(a).
Coordinate System. Our V2XPnP-Sequential dataset encompasses three coordinate systems: the LiDAR coordinate system, the camera coordinate system, and the map coordinate system. Each agent - vehicle or infrastructure - maintains its own local LiDAR and camera coordinate systems. The global map coordinate system serves as the reference for all annotations and maps. The transformation from each agent’s local LiDAR coordinate to the map coordinate in each frame is achieved with the GPS/IMU data and the offline PCD map. We also conduct the 3D-2D calibration for LiDAR and camera, as shown in Fig. 4(b).
D.3 Data Annotation and Processing
Data Annotation. The 3D bounding boxes in our V2XPnP-Sequential dataset are annotated using an open-source labeling tool, SUSTechPOINTS [23], by expert annotators. The first step is annotating the bounding boxes in the point clouds from the two CAVs and infrastructure units. Then, these bounding boxes, annotated in different agents’ coordinate frames, are processed through a V2X sequential pipeline to assign consistent object IDs across agents and temporal frames. To ensure annotation quality, each object is subjected to eight rounds of review and revision. In total, ten object categories are included in our dataset: car, pedestrian, scooter, motorcycle, bicycle, truck, van, concrete truck, bus, and road barrier. Each object annotation includes the center of the bounding box ($x,y,z$), sizes ($width,length,height$), and orientation ($roll,yaw,pitch$) in the global coordinates.
Trajectory Generation. In addition to perception data, the dataset provides a ground-truth trajectory dataset derived from the fused perception data of all agents, capturing the trajectories of objects across all frames. This trajectory dataset is primarily utilized in traditional prediction tasks, which assume all history trajectories are observable to the ego agent. However, this assumption ignores the fact that the trajectories obtained from onboard sensors are incomplete due to occlusion and limited perception range, and no specific datasets are designed to support this task. To support research in prediction with real-world sensor constraints, we provide a trajectory retrieve module in the V2XPnP-Sequential dataset to return observable trajectories of surrounding objects based on their actual visibility relationships.
Map Generation. The HD map generation involves two stages: point cloud (PCD) map generation and vector map generation. (1) To generate the PCD map, each LiDAR frame from the CAVs is preprocessed to remove dynamic objects, retaining only static elements essential for mapping. Then, a Normal Distributions Transform (NDT) scan-matching algorithm is employed to compute the relative transformation between consecutive frames, forming the basis of the LiDAR odometry. We also incorporate translation and heading information obtained from the vehicle’s GPS/IMU system, integrating them through a Kalman filter to refine the pose estimation, mitigating the drift from the error accumulation in LiDAR data. Finally, the LiDAR sequences are fused to form the PCD map across all collection areas. (2) The aggregated PCD map is imported into RoadRunner [7] to generate vector maps. Road geometry is inferred and annotated based on intensity variations visualized by distinct color mappings within RoadRunner, and all the semantic attribution is annotated based on the collected camera data, such as road type (e.g., driving, sidewalk, and parking) and line type (e.g., solid and broken yellow line combination and solid white line). Finally, the generated maps are exported in the OpenDRIVE (Xodr) format and converted to Waymo map format [9], ensuring compatibility with downstream applications.
Figure S2: Examples of intersection types in the map, including T-junctions, roundabouts, and crossroads. The gray point clouds in the background represent the PCD map, while lane transitions and gradients are depicted in the map.
D.4 Dataset Analysis
Tab. S3 presents the comparison of the V2XPnP-Sequential dataset with existing driving datasets. Our dataset tracks an average of 136 objects per scene, recording high-density and complex traffic scenarios. Furthermore, the dataset’s extensive map and trajectory data further enhance its utility in cooperative perception and prediction research across all collaboration modes. The data distribution of ten object classes is shown in Fig. 4(d). The dataset covers 24 intersections of varying types, including roundabouts, T-junctions, and crossroads, as shown in Fig. S2. Notably, many collection areas have a significant gradient, which can facilitate the detection and prediction research in diverse terrain conditions.
D.5 Dataset Privacy Protection
The V2XPnP-Sequential dataset is designed with stringent privacy safeguards to ensure the anonymity of individuals and vehicles. Trajectory data only include object IDs and positions, eliminating the possibility of tracking specific entities. All perception data undergoes privacy-preserving processing, with LiDAR annotations retaining only essential attributes such as object ID, agent type, and bounding box pose. Additionally, all image data has been anonymized, with human faces and other potentially sensitive details obscured or removed.